# Polycubes via Dual Loops

### v1

@@ Abstract
We present a complete characterization of polycubes of any genus based on their dual structure: a collection of oriented loops which run in each of the axis directions and capture polycubes via their intersection patterns. A polycube loop structure uniquely corresponds to a polycube. We also describe all combinatorially different ways to add a loop to a loop structure while maintaining its validity; similarly, we show how to identify loops that can be removed from a polycube loop structure without invalidating it. Our characterization gives rise to an iterative algorithm to construct provably valid polycube-maps for a given input surface; polycube-maps are frequently used to solve texture mapping, spline fitting, and hexahedral meshing. We showcase some results of a proof-of-concept implementation of this iterative algorithm.
@@ 1 Introduction
Polycubes are orthogonal polyhedra with axis-aligned quadrilateral faces. Intuitively one can imagine a polycube to be the result of gluing cubes together along shared faces. The simple structure of polycubes enables efficient solutions to various challenging geometric problems. Bijective mappings from general shapes to polycubes, known as polycube-maps, enable the transfer of solutions computed on polycubes to more general shapes. In this manner, polycube-maps are used to solve problems such as texture mapping [17], spline fitting [18], and hexahedral meshing [13].
@@
Formally, a polycube-map $f$ is a continuous map from a polycube $Q$ of genus $g$ to a closed $2$-dimensional surface $\mathcal{S}$ of (typically) the same genus. The edges of $Q$ map to a segmentation of $\mathcal{S}$ into patches that correspond to the faces of $Q$, known as a polycube segmentation (see Figure 1). The quality of a polycube-map is determined by two conflicting criteria: the complexity of the polycube $Q$ and the distortion introduced by the mapping $f$.
@@
Since polycubes were introduced in 2004, many methods have been proposed to construct polycube-maps for a given input surface $\mathcal{S}$, see for example [6, 8, 10]. All of these methods work in the primal: they attempt to directly create the polycube segmentation of $\mathcal{S}$ and derive the polycube and polycube-map afterwards. To do so, they use a variety of methods to construct a segmentation of $\mathcal{S}$ into surface patches and label each patch with an axis-direction $+X$, $-X$, $+Y$, $-Y$, $+Z$, or $-Z$. Such a labeled segmentation is a valid polycube segmentation if and only if a corresponding polycube exists that preserves these labels. That is, there must exist a polycube $Q$ such that $(a)$ its faces correspond one-to-one to the surface patches of $\mathcal{S}$, and $(b)$ the normal vector of each face of $Q$ corresponds to the label of its matching surface patch.
@@
Verifying whether a labeled segmentation is a valid polycube segmentation is a challenging problem. Most current primal methods rely on one of three characterization [7, 9, 19]. Eppstein and Mumford characterize simple orthogonal polyhedra [7] which are genus-0 polyhedra with simply-connected faces and exactly three mutually perpendicular axis-aligned edges meeting at every vertex. These conditions are simple to check for a segmentation, but they do not cover polycubes of higher genus or higher vertex degree (a polycube can have vertices of degree three up to six).
Figure 2: A surface segmentation (a) and the polycube it (incorrectly) maps to according to characterisations in the primal (b).
@@
The other two characterizations [9, 19] attempt to lift these restrictions. However, none of the three characterizations takes the orientation of the surface patches into account. Consider the example in Figure 2, first proposed by Sokolov [15]. The segmentation on the left has four faces parallel to the $XZ$-plane (indicated in blue), two facing up and two facing down: the top and the bottom as well as the top and the bottom of the “ramp”. According to all three characterizations, this segmentation corresponds to a polycube, namely the simple orthogonal polyhedron on the right. However, the polycube on the right clearly has three blue faces facing up and only one blue face facing down. This example shows that local conditions are not sufficient to characterize a polycube and that there are certain global constraints that must be satisfied. Methods that rely on a local characterization hence sometimes incorrectly classify labeled segmentations as valid.
There is currently no method that is guaranteed to turn an invalid labeled segmentation into a valid one, let alone do so in an efficient manner.
@@
Baumeister and Kobbelt [1] recently used a different approach: they establish a connection between quad meshes and polycubes by analyzing their dual structure. This dual structure consists of a collection of loops on the input surface $\mathcal{S}$ which run in each of the axis directions and capture global properties by virtue of their intersection patters. A version of this loop structure was described already 20 years ago by Biedl and Genc [2]. Baumeister and Kobbelt used their dual characterization as the basis for an algorithm that constructs polycube-maps for surfaces of genus 0.
Verifying that a dual loop structure corresponds to a polycube is computationally much simpler than verifying that a labeled surface segmentation corresponds to a polycube. Given a valid dual loop structure on the input surface $\mathcal{S}$ one can derive the polycube segmentation through a simple primalization step [3, 14]. In the remainder of the paper we hence focus on characterizations of polycubes via dual loops.
Both Baumeister and Kobbelt [1] and Biedl and Genc [2] characterize only polycubes of genus 0. Furthermore, neither shows how to efficiently modify an existing valid loop structure while maintaining its validity, which is a crucial requirement for any kind of optimization process [11].
@@
In this paper we describe a dual characterization of polycubes of arbitrary genus. Just as previous work, we use sets of loops labeled with the three principal axes $X$, $Y$, and $Z$. However, our loops are oriented. This seemingly small change simplifies the characterization of polycubes of genus 0 and naturally extends to polycubes of higher genus. We also describe all combinatorially different ways to add a loop to an existing loop structures while maintaining its validity. These valid loops correspond to cycles in a specific graph and can hence be detected and enumerated efficiently. Similarly we can detect all loops that can be removed from our loop structure without invalidating it.
@@
Our characterization immediately gives rise to an algorithm that constructs polycube segmentations: iteratively construct a valid dual structure on the input surface $\mathcal{S}$, starting from a simple polycube of the correct genus. At any point during the construction, there is a corresponding polycube segmentation. Add and subtract loops until the quality of the polycube segmentation is satisfactory.
@@
Our paper is organized as follows. We define polycubes and describe their properties in Section 2. In Section 3 we then present our polycube characterization via labeled and oriented loops and show in Section 4 how to modify a loop structure while maintaining validity. We implemented a proof-of-concept version of the iterative algorithm sketched above and showcase results in Section 5. Finally we conclude in Section 6 by discussing several directions for future work.
@@ 2 Polycube definition and properties
The literature on polycube-maps contains various definitions of polycubes with subtle differences. For our results, we use the fact that polycubes can be defined as a special type of quad mesh with axis-aligned faces [1]. In this case, we define the polycube as its $2$-dimensional boundary surface. We assume that the polycube has no voids, meaning the surface is connected and encloses a single bounded volume. Beyond that, this definition can be seen as a generalization of other polycube definitions.
A quadrilateral mesh (quad mesh) consists of vertices, edges, and quadrilateral faces and meets the following conditions. Each vertex is adjacent to at least one edge, and each edge is adjacent to one or two faces. Each face consists of four vertices and four edges. A quad mesh is closed if each edge is adjacent to exactly two faces. A quad mesh is orientable if a consistent circular ordering of vertices can be assigned to each face, such that edge-adjacent faces have opposite vertex orders along their common edge. A quad mesh is connected if every vertex can be reached from any other vertex by traversing edges.
Definition 1
@@
We do not impose restrictions on vertex degrees (see Figure 3a). The quadrilateral faces are not required to be unit squares (see Figure 3b), as this may not be general enough for higher genus polycubes (see Figure 3c). A polycube is also allowed to self-intersect, as this is not relevant in the context of polycube-maps [16] (see Figure 3d). We establish some basic properties of polycubes based on this definition.
Lemma 2
@@
    Any edge of the polycube connects two vertices, and by Condition 3 of Definition 1, the positions of these vertices may differ in only one coordinate. Hence, every edge must be aligned with one coordinate axis.
    Now consider a quadrilateral face $F$ of a polycube. The four interior angles of $F$ must sum up to $360^{\circ}$. As all edges are axis-aligned, each corner of $F$ must have an interior angle of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, or $270^{\circ}$. A corner with a $0^{\circ}$ angle would violate Condition 4 of Definition 1. Therefore, all corners must have an interior angle of $90^{\circ}$, making $F$ a rectangle.
Lemma 2 together with Condition 4 of Definition 1 also directly imply the following property.
Corollary 3
@@
We observe that Definition 1 makes it possible for different polycubes to have the exact same combinatorial structure, simply by changing the lengths of edges, see Figure 4. We therefore use the partial orders introduced above to define a form of combinatorial equivalence between polycubes.
Definition 4
@@
The definition above implies that a polycube $Q$ is also order-equivalent to its inverse, which can be obtained by flipping the orientation of all faces of $Q$. However, the inverse does not enclose a bounded volume (it represents a single void), and hence we will never consider it.
We will represent the partial orders $\leq_{X}$, $\leq_{Y}$, and $\leq_{Z}$ as directed graphs, namely the $X$-graph, $Y$-graph, and $Z$-graph. To avoid confusion, in the remainder of this paper we refer to the vertices in a polycube as corners and we use the term “vertex” for vertices in the $X$-, $Y$-, and $Z$-graphs.
@@ 3 Polycube dual characterization
Every embedded graph $G$ can be represented by its dual graph: a graph that has a vertex for each face of $G$, and an edge for each pair of faces of $G$ that are adjacent (share an edge). The dual of an (embedded) quad mesh is a $4$-regular graph (that is, every vertex has degree $4$), since each face in the mesh is a quadrilateral. Any $r$-regular graph, with even $r$, can be decomposed into disjoint simple cycles where consecutive edges in each cycle do not share a face [12].
We refer to an arrangement of simple closed curves $\mathcal{L}$ on a $2$-manifold as a loop structure. Such loop structures can be represented as a graph, where loop intersections define vertices, loop segments (parts of a loop between two intersections) define edges, and loop regions (regions bounded by loops) define faces (see Figure 5). As observed by Campen et al. [4], quad loop structures, the dual structures of quad meshes, are characterized by a simple set of properties.
Definition 5 ([4])
@@
Building on the works of Campen et al. [4] and Biedl and Genc [2], Baumeister and Kobbelt [1] explore how quad meshes relate to polycubes. Recognizing that polycubes are a subset of quad meshes, they examine how to transform a quad mesh into a polycube, by modifying its loop structure.
@@
In the dual loop structure of a polycube, each loop corresponds to a strip of quadrilateral faces whose center points share one coordinate [2]. These loops can be classified as $X$-, $Y$-, or $Z$-loops, see Figure 6. An $X$-loop traverses faces whose normals align with the $Y$ and $Z$ axes, that is, the normals of the faces are perpendicular to the $X$ axis. Similarly, $Y$-loops and $Z$-loops traverse faces whose normals are perpendicular to their respective axes.
The core of the dual characterization by Baumeister and Kobbelt [1] is a labeling of the loops in a loop structure with $X$, $Y$, or $Z$, resulting in a labeled loop structure. If this labeled loop structure satisfies certain properties, then the loop structure corresponds to a polycube.
@@
However, the characterization does not uniquely define a polycube. For example, the two polycubes in Figure 6 share the same labeled loop structure but differ in face alignment. The first polycube has a protruded block, while the second has an indented block. Additionally, this characterization applies only to genus-0 polycubes, and polycubes where any two dual loops intersect no more than twice. The two polycubes in Figure 6 violate both properties. Both polycubes contain a handle and are as such not genus 0. Furthermore, the rightmost $X$-loop (purple) intersects $4$ times with the middle $Y$-loop (orange). Thus, their characterization is not defined for this polycube. We address these limitations by simplifying the dual characterization of polycubes.
@@
In the remainder of this section, we fix these limitations with the following minor change to the loop structure: we assign an orientation (clockwise or counterclockwise) to every loop. The orientation of a loop can be interpreted as giving the loop two sides: a positive side where the corresponding coordinate increases in the polycube, and a negative side where the corresponding coordinate decreases. Therefore, to simplify our explanations, we orient the loops implicitly by assigning labels (positive or negative) to the two sides of a loop. This distinction between the loop sides is important for determining the direction of each polycube edge. In our figures, we use two different shades to differentiate between the two sides of each loop. The negative (--) side is represented by a lighter shade compared to the positive ($+$) side. We refer to such an extended labeled loop structure as an oriented loop structure.
Figure 7: Oriented loops and the corresponding $X$-, $Y$-, and $Z$-graphs.
@@
Before we can give a characterization of polycubes via oriented loop structures, we need to introduce a few concepts. We can use the full set of $X$-loops of an oriented loop structure to partition the underlying space (surface or polycube) into regions. We refer to these regions as $X$-zones, or simply zones in general. Similarly, we can use the $Y$-loops and $Z$-loops to obtain $Y$-zones and $Z$-zones, respectively. For each set of zones, we construct the $X$-, $Y$-, and $Z$-graph. The $X$-graph is constructed as follows (the $Y$- and $Z$-graphs are symmetric). We first add a vertex for each $X$-zone. Then, for each $X$-loop, we add a directed edge $(u,v)$, where $u$ is the zone on the -- side and $v$ is the zone on the $+$ side of the loop. We say that zones $u$ and $v$ share a loop. Since the underlying surface or polycube may have higher genus, parallel edges (where two zones share more than one loop) can occur, see Figure 7. Observe that these graphs also define the partial order of Definition 4. Consider a directed edge $(u,v)$ in the $X$-graph. Both $u$ and $v$ correspond to distinct zones which contain vertices with a shared $x$-coordinate. Notice that the shared $x$-coordinate of the vertices in zone $u$ is strictly smaller than the shared $x$-coordinate of the vertices in zone $v$.
@@
    Each zone is a surface of genus zero.
  6. 6.
    The $X$-, $Y$-, and $Z$-graph are acyclic.
In the remainder of this section we prove that, given a polycube loop structure $\mathcal{L}$ we can construct a single unique polycube $Q$ (up to order-equivalence), and conversely, that any arbitrary polycube has a corresponding dual polycube loop structure. We must first establish several key properties.
Theorem 7
@@
    Assume, for contradiction, that two loops $\lambda_{1}$ and $\lambda_{2}$ with the same axis label intersect in $\mathcal{L}$. Without loss of generality, we may assume that $\lambda_{1}$ and $\lambda_{2}$ are $X$-loops. The intersection $x$ of $\lambda_{1}$ and $\lambda_{2}$ is incident to four loop regions around $x$. Now consider the two loop regions around $x$ that are on the positive side of $\lambda_{1}$. These two loop regions are separated by $\lambda_{2}$, and hence one of these loop regions must also be on the positive side of $\lambda_{2}$. But this violates Condition 3 of Definition 6. Thus, in a polycube loop structure, loops with the same axis label cannot intersect.
Lemma 8
@@
    We only need to establish that Condition 1 of Definition 5 also holds for polycube loop structures. Assume, for the sake of contradiction, that an intersection between loops is not transversal, and let $\lambda$ be one of the loops. Then there must be a loop region incident to this intersection such that $\lambda$ occurs twice along its boundary (before and after the intersection). Clearly, both segments must have the same side label for that loop region, which violates Condition 3 of Definition 6. This is a contradiction, so every polycube loop structure is also a quad loop structure.
Theorem 9
@@
    We now verify that $\mathcal{L}$ satisfies all conditions of Definition 6:
@@
      Without loss of generality, consider an $X$-zone of $Q$. If we project this zone onto the $YZ$-plane, then it consists of a collection of disjoint simple cycles (the bounding loops) connected by interior disjoint rectangular faces whose normals are parallel to the $X$-axis. Since the projected $X$-zone is a subset of the plane, it must have genus 00. The original $X$-zone differs from the projected $X$-zone only by the (half) quadrilateral strips intersected by the bounding loops. Since adding these strips cannot increase the genus, the $X$-zone must have genus 00.
    6. 6.
@@
    Following [4], a quad mesh can be constructed from $\mathcal{L}$, where each loop region corresponds to a corner in the mesh with degree equal to the number of loop segments bounding the loop region. We now assign a position in $\mathbb{Z}^{3}$ to each corner using the $X$-, $Y$-, and $Z$-graphs of $\mathcal{L}$. Since these graphs are acyclic, we can assign integer values to each of the vertices (zones) in the graphs such that they adhere to the strict partial order induced by the graph. Because every corner $v$ is in exactly one $X$-zone, one $Y$-zone, and one $Z$-zone, we can simply assign the position $p(v)=(x,y,z)$ to $v$, where $x$, $y$, and $z$ are the integers assigned to the $X$-zone, $Y$-zone, and $Z$-zone of $v$, respectively. Let $Q$ be the resulting embedded mesh. We now show that $Q$ is a polycube using the conditions of Definition 1.

### v3

@@ Abstract
In this paper we study polycubes: orthogonal polyhedra with axis-aligned quadrilateral faces. We present a complete characterization of polycubes of any genus based on their dual structure: a collection of oriented loops which run in each of the axis directions and capture polycubes via their intersection patterns. A polycube loop structure uniquely corresponds to a polycube. We also describe all combinatorially different ways to add a loop to a loop structure while maintaining its validity. Similarly, we show how to identify loops that can be removed from a polycube loop structure without invalidating it. Our characterization gives rise to an iterative algorithm to construct provably valid polycube maps for a given input surface.
@@ 1 Introduction
Polycubes are orthogonal polyhedra with axis-aligned quadrilateral faces. The simple structure of polycubes enables efficient solutions to various challenging geometric problems. Bijective mappings from general shapes to polycubes, known as polycube maps, enable the transfer of solutions computed on polycubes to more general shapes. We distinguish between volumetric maps, which map both the interior and the surface of the polycube, and surface maps which map only the surface of the polycube to the target. Polycube maps are used to solve problems such as texture mapping [24], spline fitting [25], and hexahedral meshing [19]. We focus on surface maps only; there are established methods [10, 20, 27] that extend a surface map to a volumetric map.
@@
Formally, a polycube map $f$ is a continuous map from a polycube $Q$ of genus $g$ to a closed $2$-dimensional surface $\mathcal{S}$ of (typically) the same genus. The edges of $Q$ map to a segmentation of $\mathcal{S}$ into patches that correspond to the faces of $Q$, known as a polycube segmentation (see Figure 1). We distinguish two types of edges of $Q$: creased edges which separate two faces of different orientation and flat edges which separate two coplanar faces. In general, the quality of a polycube map is determined by two conflicting criteria: the complexity of $Q$ and the distortion introduced by the mapping $f$.
@@
Since polycubes were introduced in 2004, many methods have been proposed to construct polycube maps for a given input surface $\mathcal{S}$, see [6, 8, 9, 10, 11, 13, 14, 15, 16, 17, 26, 27, 28]. Most of these methods [6, 9, 10, 13, 14, 16, 28] work in the primal: they attempt to directly create the polycube segmentation of $\mathcal{S}$ and derive the polycube and polycube map afterwards. The resulting polycubes are represented only by their creased edges and, correspondingly, their faces can be rectilinear polygons, instead of quadrilaterals only. They use a variety of methods to construct a segmentation of $\mathcal{S}$ into surface patches and label each patch with an axis-direction $+X$, $-X$, $+Y$, $-Y$, $+Z$, or $-Z$. Such a labeled segmentation is a valid polycube segmentation if and only if a corresponding polycube exists that preserves these labels. That is, there must exist a polycube $Q$ such that $(a)$ its faces correspond one-to-one to the surface patches of $\mathcal{S}$, and $(b)$ the normal vector of each face of $Q$ corresponds to the label of its matching surface patch.
@@
Verifying whether a labeled segmentation is a valid polycube segmentation is a challenging problem. Most current primal methods rely on one of three characterizations for polycubes [7, 12, 29]. Eppstein and Mumford characterize simple orthogonal polyhedra [7] which are genus-0 polyhedra with simply-connected faces and exactly three mutually perpendicular axis-aligned edges meeting at every vertex. These conditions are simple to check for a segmentation, but they do not cover polycubes of higher genus or vertex degrees higher than three (with respect to creased edges).
Figure 2: A segmentation and the polycube it (incorrectly) maps to following the characterization of [7].
@@
The other two characterizations [12, 29] attempt to lift these restrictions. However, none of the three characterizations takes the orientation of the surface patches into account. Consider the example in Figure 2, first proposed by Sokolov [22]. The segmentation on the left has four faces parallel to the $XZ$-plane (indicated in blue), two facing up and two facing down: the top and the bottom as well as the top and the bottom of the “ramp”. According to all three characterizations, this segmentation corresponds to a polycube, namely the simple orthogonal polyhedron on the right. However, the polycube on the right clearly has three blue faces facing up and only one blue face facing down. This example shows that local conditions are not sufficient to characterize a polycube and that there are certain global constraints that must be satisfied. Methods that rely on a local characterization hence sometimes incorrectly classify labeled segmentations as valid. There is currently no method that is guaranteed to turn an invalid labeled segmentation into a valid one, let alone do so in an efficient manner.
An alternative set of primal methods [15, 26, 27] project a voxelization (the polycube) of the input surface $\mathcal{S}$ back onto $\mathcal{S}$. By construction, the polycube segmentation is directly linked to a polycube. However, the mapping might create inverted faces within the segmentation. Furthermore, the voxelization may not preserve the topology of $\mathcal{S}$. In general, voxelization methods cannot guarantee the correctness of the polycube map.
For the specific case of hexahedral meshing, there are intrinsic methods [8, 11, 17] which create volumetric polycube maps that are advantageously for downstream processing. The resulting hex meshes are generally of high quality, however, the distortion introduced by the polycube segmentation is often high and the topology of the input surface $\mathcal{S}$ may not be preserved.
@@
Baumeister and Kobbelt [1] recently used a different approach for characterizing polycubes: they represent polycubes as quad meshes, including both flat and creased edges, and analyze their dual structure. Since quad meshes have quadrilateral faces, their dual structure consists of a collection of loops which run in each of the axis directions and capture global properties by virtue of their intersection patterns. A version of this loop structure was described already 20 years ago by Biedl and Genc [2]. Both versions characterize only polycubes of genus 0. Baumeister and Kobbelt check if the collection of dual loops implied by an input quad mesh corresponds to a valid polycube according to their characterization.
Verifying that a quad mesh, including flat and creased edges, corresponds to a polycube is computationally much simpler than verifying that a labeled surface segmentation corresponds to a polycube with arbitrarily complex rectilinear faces bounded by creased edges only. Note that rectilinear faces can first be partitioned into quadrilateral faces and then the verification based on dual loops can be employed. However, partitioning a rectilinear face in such a manner that the potential validity of the corresponding polycube is maintained, is an open problem.
@@
Given an embedding of a valid dual loop structure on the input surface $\mathcal{S}$ one can derive the polycube segmentation through a simple primalization step [3, 21]. In the remainder of the paper we hence focus on the characterization and construction of (quad mesh) polycubes of arbitrary genus via dual loops. Just as previous work, we use sets of loops labeled with the three principal axes $X$, $Y$, and $Z$. However, our loops are oriented. This seemingly small change simplifies the characterization of polycubes of genus 0 and naturally extends to polycubes of higher genus.
We also describe all combinatorially different ways to add a loop to an existing loop structure while maintaining its validity. These valid loops correspond to cycles in a specific graph and can hence be detected and enumerated efficiently. Similarly we can detect all loops that can be removed from our loop structure without invalidating it.
@@
Our characterization gives rise to an algorithm that constructs polycube segmentations: iteratively construct a valid dual structure on the input surface $\mathcal{S}$, starting from a simple polycube of the correct genus. At any point during the construction, there is a corresponding polycube segmentation. We add and subtract loops until the quality of the polycube segmentation is satisfactory.
@@
Our paper is organized as follows. We define polycubes and describe their properties in Section 2. We present our polycube characterization via labeled and oriented loops in Section 3 and show how to modify a loop structure while maintaining validity in Section 4. We implemented a proof-of-concept version of the iterative algorithm sketched above and showcase results in Section 5. Finally, we conclude in Section 6 by discussing several directions for future work.
@@ 2 Polycube definition and properties
The literature on polycube maps contains various definitions of polycubes with subtle differences. For our results, we use the fact that polycubes can be defined as a special type of quad mesh with axis-aligned faces [1]. In this case, we define the polycube as its $2$-dimensional boundary surface. We assume that the polycube has no voids, meaning the surface is connected and encloses a single bounded volume. As mentioned in the introduction, we focus on surface maps between the polycube and the input surface only.
A quadrilateral mesh (quad mesh) consists of vertices, edges, and quadrilateral faces. Each vertex is adjacent to at least one edge. Each edge is adjacent to one or two faces. Each face consists of four vertices and four edges. A quad mesh is closed if each edge is adjacent to exactly two faces. A quad mesh is orientable if a consistent circular ordering of vertices can be assigned to each face, such that edge-adjacent faces have opposite vertex orders along their common edge. A quad mesh is connected if every vertex can be reached from any other vertex by traversing edges.
Definition 2.1
@@
As a consequence, our polycubes have vertices with degrees up to six (see Figure 3a). The polycube faces are not required to be unit squares (see Figure 3b), as this may not be general enough for higher genus polycubes (see Figure 3c). According to this definition, polycubes are also allowed to self-intersect, since this does not pose a problem for surface maps [23] (see Figure 3d). Note that self-intersections might cause problems for volumetric methods, such as hex meshing.
Definition 2.1 extends the most common definition of polycubes (e.g. [16]) in the sense that it allows a higher vertex degree. However, it is restricted to quadrilateral faces. Furthermore, most intrinsic methods (e.g.,[17]) allow for non-manifold vertices with overlapping edges, while our definition enforces manifold vertices.
In the following, we establish some basic properties of polycubes based on Definition 2.1.
Lemma 2.1
@@
    Any edge of the polycube connects two vertices, and by Condition 3 of Definition 2.1, the positions of these vertices may differ in only one coordinate. Hence, every edge must be aligned with one coordinate axis.
    Now consider a quadrilateral face $F$ of a polycube. The four interior angles of $F$ must sum up to $360^{\circ}$. As all edges are axis-aligned, each vertex of $F$ must have an interior angle of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, or $270^{\circ}$. A vertex with a $0^{\circ}$ angle would violate Condition 4 of Definition 2.1. Therefore, all vertices must have an interior angle of $90^{\circ}$, making $F$ a rectangle.
Lemma 2.1 together with Condition 4 of Definition 2.1 also directly imply the following property.
Corollary 2.1
@@
Definition 2.1 allows for different polycubes to have the exact same combinatorial structure, simply by changing the lengths of edges, see Figure 4. We therefore use the partial orders to define a form of combinatorial equivalence between polycubes.
Figure 4: Order-equivalence of polycubes.
Definition 2.2
@@
This definition implies that a polycube $Q$ is also order-equivalent to its inverse, which can be obtained by flipping the orientation of all faces of $Q$. However, the inverse does not enclose a bounded volume; it represents a single void. We represent the partial orders $\leq_{X}$, $\leq_{Y}$, and $\leq_{Z}$ as directed level graphs: the $X$-graph, $Y$-graph, and $Z$-graph. From now on we refer to the vertices of polycubes as corners and reserve the term vertex for vertices in the level graphs and loop structures.
@@ 3 Polycube dual characterization
Every embedded graph $G$ can be represented by its dual graph, which has a vertex for each face of $G$ and an edge for each pair of adjacent faces of $G$. The dual of an (embedded) quad mesh is a $4$-regular graph (that is, every vertex has degree $4$), since each face in the mesh is a quadrilateral. Any $r$-regular graph, with even $r$, can be decomposed into disjoint simple cycles where consecutive edges in each cycle do not share a face [18].
We refer to an arrangement of simple closed curves $\mathcal{L}$ on a $2$-manifold as a loop structure. Such loop structures can be represented as a graph, where loop intersections define vertices, loop segments (parts of a loop between two intersections) define edges, and loop regions (regions bounded by loops) define faces. Campen et al. [4] give the set of properties that characterize quad loop structures: the loop structures which are a dual of quad meshes (see Figure 5).
@@
Figure 7: Oriented loops and the corresponding level graphs. We use the colors purple, lighter purple, orange, lighter orange, green, and lighter green for $+X$, $-X$, $+Y$, $-Y$, $+Z$, $-Z$, respectively. For visual matching between the zones in the polycubes and in the level graphs, the zones are colored with two gray hues in alternating fashion.
Definition 3.1 (Campen et al. [4])
@@
Since polycubes are quad meshes they do have a dual loop structure: each loop corresponds to a strip of quadrilateral faces whose center points share one coordinate [2], see Figure 6. The loops can be classified as $X$-, $Y$-, or $Z$-loops. In our figures, we color the $X$-, $Y$-, and $Z$-loops with purple, orange, and green, respectively. An $X$-loop traverses faces whose normals align with the $Y$ and $Z$ axes, that is, the normals of the faces are perpendicular to the $X$ axis. Similarly, $Y$-loops and $Z$-loops traverse faces whose normals are perpendicular to their respective axes.
@@
Baumeister and Kobbelt [1] explore how to transform a quad mesh into a polycube, by modifying its loop structure. Specifically, they label the loop structure of an input quad mesh with $X$, $Y$, or $Z$, such that the labeled loop structure corresponds to the loop structure of a polycube. For the latter they define a characterization of the loop structure of polycubes of genus 0. However, this characterization does not uniquely define a polycube. For example, the two polycubes in Figure 6 share the same labeled loop structure but differ in face alignment. The first polycube has a protruding block, while the second has an indented block. Additionally, this characterization applies only to genus-0 polycubes, and polycubes where any two dual loops intersect no more than twice.
@@
In the remainder of this section, we fix these limitations with the following minor change to the loop structure: we assign an orientation (clockwise or counterclockwise) to every loop. The orientation of a loop can be interpreted as giving the loop two sides: a positive side where the corresponding coordinate increases in the polycube, and a negative side where the corresponding coordinate decreases. Therefore, to simplify our explanations, we orient the loops implicitly by assigning labels (positive or negative) to the two sides of a loop. This distinction between the loop sides is important for determining the direction of each polycube edge. In our figures containing oriented loops, we use two different shades to differentiate between the two sides of each loop. The negative (--) side is represented by a lighter shade compared to the positive ($+$) side. We refer to such an extended labeled loop structure as an oriented loop structure, see Figure 7.
@@
Before we can give a characterization of polycubes via oriented loop structures, we need to introduce a few concepts. We can use the full set of $X$-loops of an oriented loop structure to partition the underlying space (surface or polycube) into regions. We refer to these regions as $X$-zones, or simply zones in general. Similarly, we can use the $Y$-loops and $Z$-loops to obtain $Y$-zones and $Z$-zones, respectively. For each set of zones, we construct the $X$-, $Y$-, and $Z$-graph. The $X$-graph is constructed as follows (the $Y$- and $Z$-graphs are symmetric). We first add a vertex for each $X$-zone. Then, for each $X$-loop, we add a directed edge $(u,v)$, where $u$ is the zone on the -- side and $v$ is the zone on the $+$ side of the loop. We say that zones $u$ and $v$ share a loop. Since the underlying surface or polycube may have higher genus, parallel edges (where two zones share more than one loop) can occur, see Figure 7. Observe that these graphs also define the partial order of Definition 2.2. Consider a directed edge $(u,v)$ in the $X$-graph. Both $u$ and $v$ correspond to distinct zones which contain vertices with a shared $x$-coordinate. Notice that the shared $x$-coordinate of the vertices in zone $u$ is strictly smaller than the shared $x$-coordinate of the vertices in zone $v$.
@@
    The level graphs are acyclic.
In the remainder of this section, we prove that, given a polycube loop structure $\mathcal{L}$, we can construct a single unique polycube $Q$ (up to order-equivalence), and conversely, that any arbitrary polycube has a corresponding dual polycube loop structure. We must first establish several key properties.
Theorem 3.1
@@
    Assume, for contradiction, that two loops $\lambda_{1}$ and $\lambda_{2}$ with the same axis label intersect in $\mathcal{L}$. Without loss of generality, we may assume that $\lambda_{1}$ and $\lambda_{2}$ are $X$-loops. The intersection $x$ of $\lambda_{1}$ and $\lambda_{2}$ is incident to four loop regions around $x$. Now consider the two loop regions around $x$ that are on the positive side of $\lambda_{1}$. These two loop regions are separated by $\lambda_{2}$, and hence one of these loop regions must also be on the positive side of $\lambda_{2}$. But this violates Condition 3 of Definition 3.2. Thus, in a polycube loop structure, loops with the same axis label cannot intersect.
Lemma 3.1
@@
    We only need to establish that Condition 1 of Definition 3.1 also holds for polycube loop structures. Assume, for the sake of contradiction, that an intersection between loops is not transversal, and let $\lambda$ be one of the loops. Then there must be a loop region incident to this intersection such that $\lambda$ occurs twice along its boundary (before and after the intersection). Clearly, both segments must have the same side label for that loop region, which violates Condition 3 of Definition 3.2. This is a contradiction, so every polycube loop structure is also a quad loop structure.
Theorem 3.2
@@
    We now verify that $\mathcal{L}$ satisfies all conditions of Definition 3.2:
@@
    Thus, for every polycube $Q$ there exists a polycube loop structure $\mathcal{L}$ that forms the dual of $Q$.
@@
    Following [4], a quad mesh can be constructed from $\mathcal{L}$, where each loop region corresponds to a corner in the mesh with degree equal to the number of loop segments bounding the loop region. We now assign a position in $\mathbb{Z}^{3}$ to each corner using the level graphs of $\mathcal{L}$. Since these graphs are acyclic, we can assign integer values to each of the vertices (zones) in the graphs such that they adhere to the strict partial order induced by the graph. Because every corner $v$ is in exactly one $X$-zone, one $Y$-zone, and one $Z$-zone, we can simply assign the position $p(v)=(x,y,z)$ to $v$, where $x$, $y$, and $z$ are the integers assigned to the $X$-zone, $Y$-zone, and $Z$-zone of $v$, respectively. Let $Q$ be the resulting embedded mesh. We now show that $Q$ is a polycube using the conditions of Definition 2.1.